OpenStudy (amonoconnor):

Can someone check my answer for the following problem? Evaluate: the integral of [sin(x^(1/2))]/[x^(1/2)]•dx, with upper limits of pi^2 and (pi^2)/4 respectively.

OpenStudy (amonoconnor):

*sorry that's not the best to look at... My iPad won't let me use equations writer.

myininaya (myininaya):

\[\int\limits_{\pi^2/4 }^{\pi^2}\frac{\sin(x^\frac{1}{2})}{x^\frac{1}{2}} dx ?\]

OpenStudy (amonoconnor):

Perfect!

OpenStudy (anonymous):

\[ sin(x^1/2)/x^1/2 =sin(x^1/2) * 1/x^1/2\]So use integration by parts

myininaya (myininaya):

Substitution works just fine

OpenStudy (anonymous):

What wil u substitue?

myininaya (myininaya):

u=x^1/2 <---hint

OpenStudy (anonymous):

The top is not the derivative of the bottom

OpenStudy (anonymous):

Oh i got it

OpenStudy (anonymous):

Nice